If you miss a brick on the way up,during what remaining time interval can you reach out and catch the brick on its way down


1
Expert's answer
2021-09-20T10:00:07-0400

Explanations & Calculations


  • If you neglect (just for the ease of calculations) the air resistance (normally, air resistance can be easily neglected for a brick) it is a verticle motion under gravity.
  • Therefore, applying "\small S = ut+\frac{1}{2}at^2" for the brick's motion from the point it was missed & until it returns, the time interval can be found.

"\qquad \qquad
\begin{aligned}
\small \uparrow 0&=\small ut+\frac{1}{2}(-g)t^2\\
\small t&=\begin{cases}
\small 0:\text{not apllicable}\\\\
\large\frac{2u}{g}:\small \text{applicable}
\end{cases}
\end{aligned}"

  • Therefore depending on the speed at which it passed the catcher on its way up, this is the time available.